# CoViS-Net: A Cooperative Visual Spatial Foundation Model for Multi-Robot Applications

### v1

@@ Abstract
Spatial understanding from vision is crucial for robots operating in unstructured environments. In the real world, spatial understanding is often an ill-posed problem. There are a number of powerful classical methods that accurately regress relative pose, however, these approaches often lack the ability to leverage data-derived priors to resolve ambiguities. In multi-robot systems, these challenges are exacerbated by the need for accurate and frequent position estimates of cooperating agents. To this end, we propose CoViS-Net, a cooperative, multi-robot, visual spatial foundation model that learns spatial priors from data. Unlike prior work evaluated primarily on offline datasets, we design our model specifically for online evaluation and real-world deployment on cooperative robots. Our model is completely decentralized, platform agnostic, executable in real-time using onboard compute, and does not require existing network infrastructure. In this work, we focus on relative pose estimation and local Bird’s Eye View (BEV) prediction tasks. Unlike classical approaches, we show that our model can accurately predict relative poses without requiring camera overlap, and predict BEVs of regions not visible to the ego-agent. We demonstrate our model on a multi-robot formation control task outside the confines of the laboratory.
I Introduction
Spatial understanding is a cornerstone of robotic operation in unstructured environments, relying on effective pose estimation and environmental perception [48, 46, 8, 58, 49, 7, 28, 29]. While traditional methods leveraging GNSS, LiDAR, Time of Flight (ToF), and Ultra-Wide Band (UWB) sensors have been instrumental in advancing robotics, they often come with constraints (e.g., indoor, outdoor, sunny, IR-reflective surfaces, etc.). Unlike these methods, color cameras offer a low-cost, energy-efficient, and rich data source suitable for many environments, aligned with the vision-centric design principles of real-world, human-designed environments.
Classical vision-based techniques such as Visual Odometry [48], Visual SLAM [7], and Structure-from-Motion [49], struggle with the inherent ill-posed nature of spatial tasks, lacking the capability to integrate human-like semantic priors to resolve ambiguous situations. These challenges are exacerbated in multi-robot applications, requiring not only environmental spatial understanding but also rapid and accurate relative pose estimates to other robots, as seen in scenarios such as flocking [39], path planning [50], and collaborative perception [69]. Explicit detection of other agents is a viable alternative [58, 8, 46, 28, 29], but requires line-of-sight measurements and is not platform agnostic. On the other hand, deep pose predictors demonstrate the ability to learn human priors to solve complex spatial tasks [27, 37, 63, 64, 6, 3, 40]. Unfortunately, such methods have yet to be broadly integrated and demonstrated on robots in real-world scenarios – particularly when operating under hardware or computational constraints. When considering multi-robot systems, prior assumptions on data availability and latency are broken, and otherwise promising work fails to deliver. Furthermore, uncertainty metrics are essential to reason about the reliability of predictions.
@@ I Introduction
Figure 1: Our model can be used to control the relative pose (yellow and purple circle) of multiple follower robots (yellow and purple cone) to a leader robot (with blue cone) following a reference trajectory using visual cues in the environment only. Even if there is no direct visual overlap (cone intersection), the model is able to correctly estimate the relative pose.
@@
Acknowledging these limitations, our work introduces CoViS-Net, a decentralized, platform-agnostic visual spatial foundation model designed for real-world multi-robot applications. This model operates in real-time on onboard computers without the need for pre-existing static network infrastructure, thus addressing critical scalability, deployment flexibility, and operational robustness challenges.
We design a deep distributed stateless architecture to focus on the challenges specific to multi-robot systems, including the need for relative pose estimation [39, 50, 56, 26, 67] and the prediction of BEVs of regions without local camera coverage [21, 43, 62, 22]. We demonstrate the efficacy of our approach through real-world applications, showcasing its use on a multi-robot control task outside laboratory settings.
In summary, our contributions are threefold:
  1. 1.
@@
    A novel architecture for platform-agnostic, decentralized, real-time, multi-robot pose estimation from monocular images, incorporating uncertainty awareness without the need for a predefined map.
  2. 2.
    Extension of this architecture to predict uncertainty-aware BEVs, enhancing spatial understanding in multi-robot systems, especially in occluded or unobserved areas.
  3. 3.
    Real-world validation of our model’s applicability to multi-robot control and BEV estimation, even for ill-posed situations where images share no overlap.
II Related Work
Multi-Robot Systems: Multi-robot systems necessitate the coordination of multiple, often fast-moving robots within close proximity in real-world settings. This coordination requires not only a spatial understanding of the immediate environment but also accurate and quick pose estimates of cooperating robots. Key applications include flocking, where agents maintain proximity using rules of separation, alignment, and cohesion [39]; formation control, where agents maintain or adjust formations to navigate obstacles [31]; and trajectory deconfliction to prevent agent collisions (RVO) [56]. Additional scenarios involve multi-agent object manipulation [26, 2] and area coverage [67, 51]. Recent advances employ Graph Neural Networks (GNNs) for enhanced multi-agent control and cooperative perception, facilitating inter-robot communication through latent messages to address partial knowledge challenges [55, 14, 15, 5, 4, 59, 69, 21, 43, 62, 22]. A significant gap in this research area is the acquisition and reliance on pose knowledge, typically relative pose, which we address in this paper. For learning-based approaches, our model can be used as a plug-in model to fine-tune for more specific use cases.
Map-Free Relative Pose Regression: Map-Based Pose Regression, as in [18, 45], predicts a camera image’s pose within a learned model of a scene. Conversely, Map-Free Relative Pose Regression, exemplified by [3], estimates the relative pose between two camera images without needing scene-specific training. Feature matching between images retrieves orientation and a scale-less transformation vector. Traditional non-learning methods like SIFT [25] handle images with significant overlap and minimal differences effectively. However, learning-based approaches for feature matching, such as Superglue [44], LOFTR [53], Gluestick [33], and Lightglue [23], have gained popularity, even though they still require considerable visual overlap. Integrating scale is possible by pairing these methods with depth maps or directly through Siamese network architectures that use dual CNNs for image encoding and pose estimation via a fully connected layer [27, 20, 12, 1, 63, 37, 64]. These models, generally assuming visual overlap and using mean-squared-error loss, lack uncertainty estimates, limiting their applicability to robotics. Robotics applications of similar methods [63, 37, 64] have not been tested in real deployments. To the best of our knowledge, Winkelbauer et al. [61] is the only method to employ uncertainty estimation in robot localization through Monte Carlo dropout, which is an epistemic uncertainty. In this work, we introduce a method to predict aleatoric uncertainties. Extending beyond field-of-view pose prediction, [6] and [3] proposed novel rotation and position estimation methods in challenging scenarios. Learning-based limitations are highlighted by [68], advocating for essential matrix estimation, and [40], using vision transformers for direct essential matrix estimation. Our approach diverges from this prior body of work by providing reliable relative poses with uncertainty estimates, without requiring visual overlap, and using unmodified pre-trained foundation models, enabling transfer learning [60].
BEV Prediction: In autonomous vehicle applications, detailed environmental representation is crucial. The BEV, merging data from diverse sensors in an agent-centric frame, is particularly effective. Traditional BEV generation involves multiple steps, but deep-learning approaches have shown promise. Pioneers like Roddick and Cipolla [41] projected image features into a BEV representation for prediction. The Lift, Splat, Shoot methodology by Philion and Fidler [36] adopted a novel multi-camera rig approach. BEVFormer [65], PETR [24], and BEVSegformer [34] contributed new dimensions, such as spatiotemporal information and multi-task learning. Extensions to multi-agent scenarios [62, 43, 21, 22, 59, 16] predominantly relied on GNSS for global pose knowledge. Inspired by Dutta et al. [11], our work employs a transformer-based BEV predictor, mapping transformer output sequences directly to a BEV grid. We focus on a 6m x 6m indoor area, unlike the larger scales typical in autonomous driving. Beyond [11], we integrate pose embeddings from our estimator into multi-node feature aggregation, showing how the infusion of explicit pose information provides benefits over simple aggregation.
III Problem Formulation
Given a multi-robot system, our goal is (i) for each robot to predict its pose with respect to other robots based only on non-line-of-sight visual correspondences in the environment, and (ii) to use this information for downstream tasks, such as for estimating a local occupancy grid map or for real-time multi-robot control in transfer learning [60].
Consider a multi-robot system represented by a set of nodes $\mathcal{V}$. Each node $v_{i}\in\mathcal{V}$ has a position $\mathbf{p}_{w,i}$ and an orientation $\mathbf{R}_{w,i}$ in the world coordinate frame $\mathcal{F}_{w}$. The set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, defined by the communication topology, connects nodes $v_{i}$ to $v_{j}$ and defines the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. For each edge $(v_{i},v_{j})\in\mathcal{E}$, the relative position of $v_{j}$ in the coordinate frame of $v_{i}$ (denoted as $\mathcal{F}_{i}$) is $\mathbf{p}_{i,ij}=\mathbf{R}_{w,i}^{-1}\cdot(\mathbf{p}_{w,j}-\mathbf{p}_{w,i})$, and the relative rotation is $\mathbf{R}_{i,ij}=\mathbf{R}_{w,i}^{-1}\cdot\mathbf{R}_{w,j}$.
Each node $v_{i}$ is equipped with a camera $C_{i}$, located and oriented identically to the robot. The camera output at node $v_{i}$ is represented as an image $I_{i}$. From the images, we must estimate relative poses and associated uncertainties ${(\hat{\mathbf{p}}_{i,ij,,}\>\hat{\mathbf{R}}_{i,ij},\>\sigma^{2}_{ij})\>\forall\>v_{j}\in\mathcal{N}(v_{i})}$ with respect to $v_{i}$ in its coordinate frame $\mathcal{F}_{i}$. Notably, we do not assume that there is overlap between images, nor line-of-sight positioning between robots. Relying exclusively on common image features makes this approach platform-agnostic and only requires a single monocular camera. Utilizing these estimated poses, the robots are tasked to cooperatively generate a BEV $\hat{\mathrm{BEV}_{i}}$ around each $v_{i}$. The pose estimates and BEV can be used as input for a wide range of downstream multi-robot applications.
IV Model Training
In this section, we explain the whole training pipeline of our foundation model. We generate a simulation dataset and define our model architecture before we introduce our uncertainty estimation strategy and training protocol.
@@ IV Model Training
Figure 2: We visualize a sample from the simulation validation dataset $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ for five nodes, corresponding to the five columns. Each node is associated with the colored border in the first row. The top row shows the camera image $I_{i}$ from the perspective of each node. The middle row shows the five ground truth labels for each respective node in the coordinate system $\mathcal{F}_{i}$. Each robot is centered in its coordinate frame, facing upwards at position (0, 0). The ground truth poses $\mathbf{p}_{i,ij}$ of the other robots are indicated, together with their field of view. The background shows the BEV map $\mathrm{BEV}_{i}$ for each robot. The bottom row shows the corresponding predictions for each robot. We only show pose predictions $\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij}$ with a sufficiently low uncertainty $\sigma^{2}_{i}{ij}$. In the background, the predicted corresponding BEV $\hat{\mathrm{BEV}_{i}}$ is shown.
IV-A Dataset
We utilize the Habitat simulator [47, 54] along with the training set of the HM3D dataset [38], which comprises $800$ scenes of 3D-scanned real-world multi-floor buildings. We select Habitat and HM3D for their photorealism and the availability of ground-truth pose information, as well as detailed navigation mesh information that can be converted to a BEV of the scene. We instantiate a single agent equipped with a calibrated camera with a field of view of $120^{\circ}$. For each scene, we randomly sample several uniform agent positions based on the total area of the scene. For each sample, the camera orientation is randomized across all three axes. Additionally, for each scene, we extract the processed navigation mesh, which indicates areas that are navigable by a mobile robot, or not navigable due to obstacles or overhangs. In total, we extract $3,816,288$ images from $1,411$ floors in $800$ scenes.
We divide this data into training $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ ($80\%$), testing $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ ($19\%$), and validation $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ ($1\%$) sets. From these sets, we sample pairs of $N_{\mathrm{max}}=5$ images and poses by (1) uniformly sampling $25\%$ of individual images from each dataset, and then (2) uniformly sampling $N_{\mathrm{max}}-1$ images within a distance of $d_{\mathrm{max}}=2\textrm{ m}$ around each sample from (1). This process yields $763,257$ training and $276,680$ test tuples, each of which are at most $2d_{\mathrm{max}}$ apart and pointing in random directions, thus providing realistically close and distant data samples for estimating both poses and uncertainties in case of no overlaps.
IV-B Model
@@
Figure 3: Overview of the model architecture: We illustrate the decentralized evaluation with respect to $v_{i}$, which receives embeddings $\mathbf{E}_{i},\mathbf{E}_{j}$ generated by the encoder $f_{\textrm{enc}}$ from the neighbors $v_{j},v_{k}\in\mathcal{N}(v_{i})$. For each received embedding, we employ $f_{\textrm{pose}}$ to compute the pose. This pose can then be utilized for pose control or any other relevant task. We concatenate the pose embedding $\mathbf{E}_{ij}$ with the image embedding $\mathbf{E}_{i}$ and subsequently aggregate this into a node feature $\mathbf{F}_{i}$.

### v2

@@ Abstract
Autonomous robot operation in unstructured environments is often underpinned by spatial understanding through vision. Systems composed of multiple concurrently operating robots additionally require access to frequent, accurate and reliable pose estimates. Classical vision-based methods to regress relative pose are commonly computationally expensive (precluding real-time applications), and often lack data-derived priors for resolving ambiguities. In this work, we propose CoViS-Net, a cooperative, multi-robot visual spatial foundation model that learns spatial priors from data, enabling pose estimation as well as general spatial comprehension. Our model is fully decentralized, platform-agnostic, executable in real-time using onboard compute, and does not require existing networking infrastructure. CoViS-Net provides relative pose estimates and a local bird’s-eye-view (BEV) representation, even without camera overlap between robots, and can predict BEV representations of unseen regions. We demonstrate its use in a multi-robot formation control task across various real-world settings. We provide supplementary material online [^1] https://sites.google.com/view/covis-net, [^2] We will open source our trained model in due course..
Keywords: Multi-Robot Systems, Robot Perception, Foundation Models
@@
Figure 1: Our model can be used to control the relative pose of multiple follower robots (depicted in yellow and magenta) to a leader robot (blue) following a reference trajectory using visual cues in the environment only (the field-of-view is depicted by cones). Even if there is no direct visual overlap (cone intersection), the model is able to estimate the relative pose.
1 Introduction
Spatial understanding is a cornerstone of robotic operation in unstructured environments, relying on effective pose estimation and environmental perception [1, 2, 3, 4, 5, 6, 7, 8]. While GNSS, LiDAR, and UWB sensors have been instrumental in advancing robotics, they are limited by constraints such as indoor operation, unreliability when operating around reflective surfaces or in bright environments. In contrast, color cameras offer a low-cost, energy-efficient, and rich data source, aligned with the vision-centric design of real-world, human-designed environments. Classical vision-based techniques [1, 6, 5] struggle with the ill-posed nature of image processing and cannot leverage semantic priors to resolve ambiguous situations. These challenges are exacerbated in multi-robot scenarios that require fast and accurate relative pose estimates for tasks like flocking [9], path planning [10], and collaborative perception [11]. Explicit agent detection [4, 3, 2, 7, 8] requires line-of-sight and is not platform-agnostic, whereas deep pose predictors [12, 13, 14, 15, 16, 17, 18] show promise but have are not widely adopted in multi-robot systems due to hardware constraints. Additionally, incorporating uncertainty metrics is essential for assessing the reliability of predictions in these applications.
@@ 1 Introduction
Acknowledging these limitations, we introduce CoViS-Net, a decentralized, platform-agnostic visual spatial foundation model designed for real-world multi-robot applications. This model operates in real-time on onboard computers without the need for pre-existing network infrastructure, thus addressing critical scalability, flexibility, and robustness challenges. We design a deep distributed stateless architecture to tackle multi-robot uncertainty-aware relative pose estimation and BEV representation prediction for regions without local camera coverage. We demonstrate our approach in real-world applications, including a multi-robot control task outside of the laboratory.
@@
In summary, our contributions are threefold: (1) We introduce a novel architecture for platform-agnostic, decentralized, real-time multi-robot pose estimation from monocular images, incorporating uncertainty awareness without a predefined map. (2) We extend it to predict uncertainty-aware BEV representations, facilitating spatial understanding in occluded areas. (3) We validate our model in the real-world for a multi-robot control task, and show its capabilities even in ill-posed situations where images share no overlap.
@@
Figure 2: By leveraging priors, such as the way humans lay out their rooms, our model provides rough relative positions even without overlapping fields of view, which classical approaches cannot achieve. This figure shows a sample from the simulation validation dataset $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ for five nodes, each represented by a column. The top row displays the camera image $I_{i}$ from each node’s perspective. The middle row shows the ground truth labels in the coordinate system $\mathcal{F}_{i}$, with each robot centered and facing upwards at (0, 0), and the ground truth poses $\mathbf{p}_{i,ij}$ of other robots indicated, along with their field of view. The background shows the BEV representation $\mathrm{BEV}_{i}$ for each robot. The bottom row presents the corresponding predictions, displaying pose predictions $\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij}$ with low uncertainty $\sigma^{2}_{\mathrm{p},ij}\>\sigma^{2}_{\mathrm{R},ij}$, predicted uncertainties $\sigma^{2}_{\mathrm{p},i}{ij}$ as dashed oval and $\sigma^{2}_{\mathrm{R},i}{ij}$ similar to the FOV, and the predicted BEV representation $\hat{\mathrm{BEV}}_{i}$ in the background.
2 Related Work
Multi-Robot Systems: Multi-robot systems require coordination of fast-moving robots in close proximity, necessitating spatial understanding and accurate pose estimates. Key applications include flocking [9], formation control [19], trajectory deconfliction [20], multi-agent object manipulation [21, 22], and area coverage [23, 24]. Recent advances use Graph Neural Networks (GNNs) for enhanced control and perception, enabling inter-robot communication through latent messages to address the challenge of partial observability [25, 26, 27, 28, 29, 30, 11, 31, 32, 33, 34]. However, a significant gap remains in acquiring relative pose knowledge with uncertainty estimates, which our model addresses.
Map-Free Relative Pose Regression: Map-Free Relative Pose Regression [17] estimates the relative pose between two camera images without scene-specific training. Traditional methods like SIFT [35] and learning-based approaches such as Superglue [36], LOFTR [37], Gluestick [38], and Lightglue [39] require significant visual overlap between views. Integrating scale involves depth maps or Siamese networks [12, 40, 41, 42, 14, 13, 15]. These models lack uncertainty estimates, limiting their robotics applicability, and have not been tested in real-world, real-time deployments. Our approach provides reliable relative poses with aleatoric uncertainty estimates, beyond visual overlap constraints, using unmodified pre-trained foundation models, enabling transfer learning [43].
BEV Representation: The construction of BEV representations is an effective way of providing detailed spatial comprehension by merging data from a set of sensors. Techniques like Pyramid Occupancy Networks [44] and Lift, Splat, Shoot [45] project image features into a BEV representation. BEVFormer [46], PETR [47], and BEVSegformer [48] add dimensions such as spatiotemporal information and multi-task learning. Multi-agent scenarios [33, 32, 31, 34, 30, 49] rely on GNSS for global pose knowledge, precluding application to indoor environments. Our work proposes a transformer-based BEV representation predictor that utilizes relative poses predicted from monocular images, instead of known static poses of a camera rig or poses derived from external (infrastructural) sensors.
3 Problem Formulation
Given a multi-robot system, our goals are: (i) for each robot to predict its pose and uncertainty relative to other robots as well as corresponding embeddings using non-line-of-sight visual correspondences, and (ii) to use these embeddings for downstream tasks like local occupancy grid mapping or real-time multi-robot control in transfer learning [43]. Consider a multi-robot system represented by a set of nodes $\mathcal{V}$. Each node $v_{i}\in\mathcal{V}$ has a position $\mathbf{p}_{w,i}$ and an orientation $\mathbf{R}_{w,i}$ in the world coordinate frame $\mathcal{F}_{w}$. The set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ defines the communication topology, connecting nodes $v_{i}$ to $v_{j}$ and defining the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. For each edge $(v_{i},v_{j})\in\mathcal{E}$, the relative position of $v_{j}$ in the coordinate frame of $v_{i}$ (denoted as $\mathcal{F}_{i}$) is $\mathbf{p}_{i,ij}=\mathbf{R}_{w,i}^{-1}\cdot(\mathbf{p}_{w,j}-\mathbf{p}_{w,i})$, and the relative rotation is $\mathbf{R}_{i,ij}=\mathbf{R}_{w,i}^{-1}\cdot\mathbf{R}_{w,j}$. Each node $v_{i}$ is equipped with a camera $C_{i}$. The camera output at node $v_{i}$ is represented as an image $I_{i}$. From the images, we estimate relative poses and associated uncertainties ${(\hat{\mathbf{p}}_{i,ij},\>\sigma^{2}_{\mathrm{p},ij},\>\hat{\mathbf{R}}_{i,ij},\>\sigma^{2}_{\mathrm{R},ij})\>\forall\>v_{j}\in\mathcal{N}(v_{i})}$ with respect to $v_{i}$ in its coordinate frame $\mathcal{F}_{i}$. This approach does not assume overlap between images or line-of-sight visibility between robots, relying exclusively on the relationship between image features. This makes it platform-agnostic and requires only a single monocular camera. Using these estimated poses, the robots cooperatively generate a bird’s eye view representation $\hat{\mathrm{BEV}}_{i}$ around each node $v_{i}$, which can be used for various downstream multi-robot applications.
4 Approach
In this section, we elaborate the architecture and training pipeline of our foundation model. We detail our simulation dataset and define our model architecture, before introducing the uncertainty estimation strategy and training protocol.
4.1 Model
@@
Figure 3: Overview of the model architecture: We illustrate the decentralized evaluation with respect to ego-robot $v_{i}$, which receives embeddings $\mathbf{E}_{j},\mathbf{E}_{k}$ generated by the encoder $f_{\textrm{enc}}$ from the neighbors $v_{j},v_{k}\in\mathcal{N}(v_{i})$. For each received embedding, we employ $f_{\textrm{pose}}$ to compute the pose. We concatenate the pose embedding $\mathbf{E}_{ij}$ with the image embedding $\mathbf{E}_{j}$ (and $\mathbf{E}_{ik}$ with $\mathbf{E}_{k}$) and subsequently aggregate this into a node feature $\mathbf{F}_{i}$. Model outputs are framed with a dashed line.